OpenStudy (anonymous):

A piece of wire 25 m long is cut into two pieces. One piece is bent into a square and the other is bent into a circle. (a) How much of the wire should go to the square to maximize the total area enclosed by both figures? ANS= 0 m *** (b) How much of the wire should go to the square to minimize the total area enclosed by both figures?

OpenStudy (anonymous):

if you want a guess i would say put everything in to the circle and forget about the square

OpenStudy (anonymous):

oh sorry you already answered that

OpenStudy (anonymous):

probably all in the square to minimize rigth? just a guess though

OpenStudy (anonymous):

that's my guess too, but, i'll bring mathematica out again ;-)

OpenStudy (anonymous):

yeah, mathematica says to put 100% into the square

OpenStudy (anonymous):

In[940]:= Minimize[{\[Pi]r^2 + x^2, 4 x + 2 \[Pi]r == 25}, {r, x}] Out[940]= {1/16 (625 - 100 \[Pi]r + 20 \[Pi]r^2), {r -> 0, x -> 1/4 (25 - 2 \[Pi]r)}}

OpenStudy (anonymous):

How much of the wire should go to the square to minimize the total area enclosed by both figures? i think it requires a numerical answer

OpenStudy (anonymous):

all of it. 25 m

OpenStudy (anonymous):

it says it's incorrect :S

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

oh now i get x=3.5, that's one side of the square

OpenStudy (anonymous):

sorrry, still incorrect T___T;

OpenStudy (anonymous):

well it's 4 times that i just gave you one side

OpenStudy (anonymous):

ohh, i did the area (3.5)^2 instead of (3.5)*4. thanksss

OpenStudy (anonymous):

did it say that was right?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

(b) How much of the wire should go to the square to minimize the total area enclosed by both figures? The exact answer is:\[\frac{100}{4+\pi } \text{ meters}\]or 14.002479 meters to 8 digits.

OpenStudy (anonymous):

Take the derivative of the expression of interest\[D\left[\left(\frac{x}{4}\right)^2+\frac{(25-x)^2}{4 \pi },x\right] \]set it to zero\[-\frac{25-x}{2 \pi }+\frac{x}{8}==0 \]and solve for x.\[x\to \frac{100}{4+\pi } \]A plot is attached.

OpenStudy (anonymous):

Using Mathematica's Minimize function:\[\text{Minimize}\left[ \left\{\left(\frac{x}{4}\right)^2+\frac{(25-x)^2}{4 \pi }\right\},x\right]\to \]\[\left\{\frac{2500+\frac{40000}{(4+\pi )^2}+\frac{10000 \pi }{(4+\pi )^2}-\frac{20000}{4+\pi }}{16 \pi },\left\{x\to \frac{100}{4+\pi }\right\}\right\}\text{ //N} \]